Problem 1- Saving Money You want to deposit money into a bank account that pays 1.5% compounded monthly. Answer the following questions: f depositing $4000, how much is avaiable in the account at the end of S years? .How much do you have to deposit now in order to have $5000 available at the end of 5 years? Instead of depositing one sum of money at the beginning of the 5-year period, suppose you wish to make monthly payments. How much needs to be deposited at the end of each month in order to have $5000 available at the end of the 5 years? Also, how much interest is earned over the 5 years?
Answer #1

Answer.

Part 1  

T= 60 months; R= 1.5%p.m; principle=$4000

F.V = P.V*e^rt

=$4000*2.71828^1.5%*60

=$4000*2.71828^.9

=$4000*2.45960311116

=$9838.4

Part 2

T= 60 months; R= 1.5%p.m; principle=?

P.V= F.V/e^r*t

P.V= $5000/2.71828^.9

=$5000/2.4596031116

=$2032.84.

part 3

assume amount to be invested monthly be X

F.V=p.v*(1+i)^n

$5000=X*(1+1.5%)^60+X*(1+1.5%)^59+................X*(1+1.5%)^1)+X

$5000=X(1+1.5%)^60+(1+1.5%)^59+...........(1+1.5%)^1)

$5000=X(2.44+2.40+2.37++2.33+2.30+2.27+2.24+.........1.015)

X=$5000/(2.44+2.40+2.37++2.33+2.30+2.27+2.24+.........1.015)

You can get the answer by doing above step. i do not have at this time calculator to solve this ..